In a 5.00 km race, one runner runs at a steady 12.0 km/h and another runs at 14.0 km/h .
1
Expert's answer
2019-12-20T10:25:30-0500

How long does the faster runner have to wait at the finish line to see the slower runner cross?


"t=\frac{d}{v_1}-\frac{d}{v_2}"

"t=\frac{5}{12}-\frac{5}{14}=0.06\ h=3.6\ min"